# Advancing Human Action Recognition with Foundation Models trained on Unlabeled Public Videos

### v3

@@ Abstract
The increasing variety and quantity of tagged multimedia content on a variety of online platforms offers a unique opportunity to advance the field of human action recognition. In this study, we utilize 283,582 unique, unlabeled TikTok video clips, categorized into 386 hashtags, to train a domain-specific foundation model for action recognition. We employ VideoMAEv2, an advanced model integrating Masked Autoencoders (MAE) with Vision Transformers (ViT), pre-trained on this diverse collection of unstructured videos. Our model, fine-tuned on established action recognition benchmarks such as UCF101 and HMDB51, achieves results comparable with the state-of-the-art: 97.5% on UCF101, 83.9% on HMDB51, 81.1% on Kinetics-400, and 69.2% on Something-Something V2 using the ViT-base backbone. These results highlight the potential of using TikTok videos as a valuable source of diverse and dynamic content for training foundation models. Our investigation confirms that while initial increases in pre-training data volume significantly enhance model performance, the gains diminish as the dataset size continues to expand. Our findings identify two critical lessons in self-supervised learning for computer vision: (1) additional pre-training data yields diminishing benefits and (2) these diminishing benefits can be observed at relatively small dataset sizes on the order of thousands of images.
@@ 1 Introduction
Action recognition is a fundamental computer vision task that involves identifying and classifying human actions in video sequences, with significant applications in areas such as security, interactive gaming, and healthcare. In the era of social media platforms, TikTok has emerged as a significant source of human action videos. We utilized a collection of 283,582 TikTok video clips, encompassing a wide range of human activities, to train a foundation model tailored to human action recognition. This internally curated collection aims to represent various human actions relevant to real-world scenarios and cultural contexts. Importantly, it includes well-defined but non-standard human actions, addressing gaps in existing action recognition datasets and diversifying the range of human activities recognized by current models.
@@
Unlike Kinetics [4], which also includes user-generated content from YouTube, our collection exclusively features TikTok videos, known for their dynamic and spontaneous nature. This distinction offers unique training advantages, as such videos often exhibit quick transitions, novel perspectives, and creative interactions—elements less prevalent in YouTube’s often longer-form content. These characteristics provide robust challenges for training models, enhancing their ability to interpret a wide range of human actions in diverse and unpredictable environments.
@@
We conducted a series of experiments to assess the potential of videos collected in this manner for pre-training computer vision foundation models for action-related tasks. We employed the well-established, state-of-the-art model VideoMAE V2, using it with ViT-base backbones[9]. The ViT-base was used to provide a baseline comparison with other models using similarly sized backbones. Subsequent fine-tuning on well-known video action recognition datasets like UCF101[7], HMDB51[5], Kinetics[4], and Something-Something V2[3] demonstrated valuable insights: the performance achieved with both backbones was comparable to, or in some cases even better than, models trained exclusively on these conventional datasets or even those pre-trained on much larger datasets, where performance increases are often marginal.
We also explored the relationship between the quantity of pre-training videos and the performance on downstream tasks. By sampling subsets of our collection, starting with 1,000 videos and increasing in increments of 1,000 up to 6,000 videos, and pre-training VideoMAE V2[9] on these subsets, we observed that performance remains robust as long as the number of pre-training videos stays above a certain threshold, highlighting that massive datasets may not always be necessary for effective pre-training. This exploration not only emphasizes the efficiency of using online unlabeled videos as a pre-training source but also challenges the conventional reliance on extensive video datasets for achieving high model performance. Before deployment of this approach in the real world, we encourage researchers to think through the ethics of data scraping from publicly available sources to train deep learning models.
@@ 2.1 Human Action Recognition Datasets
The evolution of action recognition research has been strongly influenced by the availability of benchmark datasets. The UCF101[7] dataset, with its diverse collection of human actions, has been a cornerstone for early research in this domain, guiding and benchmarking numerous computer vision innovations applied to action recognition. Similarly, the HMDB-51[5] dataset, despite its limited size, has been pivotal in providing a varied set of human actions captured under realistic conditions.
The Kinetics Human Action Video dataset[4], introduced in 2017, marked a significant milestone in this field. With its vast collection of YouTube-sourced videos encompassing 400 distinct human action classes, Kinetics has become an indispensable resource for researchers, setting a higher benchmark and promoting advanced research in action recognition.
Another noteworthy dataset is the Something-Something[3] collection. Its unique focus on human-object interactions captured in real-world scenarios has offered researchers a different perspective, emphasizing the complexity and diversity of everyday actions.
These datasets, along with several others, have collectively laid the foundation for the current advancements in action recognition, facilitating the development and evaluation of state-of-the-art models and techniques. Recognizing the rapid evolution of content on platforms like TikTok and the emergence of diverse, contemporary actions and scenarios, we are motivated to develop a comprehensive dataset highlighting the latest trends and user behaviors specific to new short-video social media platforms. Our aim was to provide the community with a dataset with well-defined yet non-standard human actions that complement common actions already represented in current action recognition datasets, thus adding variety to the actions that action recognition models are currently evaluated for.
@@ 2.2 Computer Vision Foundation Models
Foundation models have become essential tools in the field of computer vision. The advent of models like Florence[10] showcases the versatility and scope of these models. Specifically, Florence encapsulates the progression of foundation models, adeptly offering transitions from broad scene outlines to detailed object-centric views and from static imagery to dynamic video sequences. Additionally, its effectiveness spans various modalities, from RGB imagery to textual data, establishing new standards in tasks like classification, object detection, retrieval, and particularly in action recognition.
@@
Another significant model in this domain is VideoMAE V2[9]. Introduced at CVPR 2023, this self-supervised pre-training methodology stands out for its ability to train on datasets exceeding 1 million videos. With its innovative dual masking strategy and progressive training approach, VideoMAE V2 sets new standards in action recognition. Empirical evaluations demonstrate its superior performance across multiple benchmarks, emphasizing its effectiveness in real-world scenarios. We anticipate that our extensive TikTok dataset will serve as a valuable resource for the continued development and refinement of computer vision foundation models, contributing to the advancement of this rapidly evolving field.
@@ 4 Performance on Benchmarks
In the domain of human action recognition, benchmarking performance is essential to understand the efficacy of a model. This section evaluates the performance of our model, initially pre-trained on the entire collection of 283,582 unlabeled TikTok video clips using VideoMAE V2[9], and subsequently fine-tuned on Kinetics-400 (K400). Following this, the model underwent further fine-tuning on UCF-101[7] and HMDB51[5], adopting a transfer learning approach akin to the experimental setups used in VideoMAE V2 studies. Additionally, we extended our evaluation to include the Something-Something V2[3] dataset, assessing the model’s adaptability and performance across varied action recognition contexts.
@@ 4.1 Pre-training on VideoMAE V2
VideoMAE V2[9] is a state-of-the-art masked autoencoding framework, notable for its flexibility in backbone selection, ranging from base to billion-level configurations. We opted for the ViT-base backbone, which, while less computationally intense than its billion-level counterparts, significantly enhances performance across various video-related tasks, including video action recognition. The generalization capability of VideoMAE V2 pre-trained ViTs[1] as video foundation models has been demonstrated across multiple downstream tasks, highlighting its robustness and adaptability.
Our training was conducted on 8 NVIDIA A100 40GB GPUs, utilizing the full set of 283,582 unlabeled TikTok videos. The pre-training on the ViT-base backbone lasted for 20 days, providing a comprehensive base for developing a robust model capable of understanding a wide array of human actions.
@@ 4.2 Fine-Tuning on Established Datasets
After initial pre-training, the model underwent two distinct paths of fine-tuning. The first path involved direct fine-tuning on the Kinetics-400 and Something-Something V2 datasets to assess immediate applicability to broad action recognition tasks. The results of this direct fine-tuning are detailed in Table 2. The second path involved using the model fine-tuned on Kinetics-400 as an intermediary for transfer learning on UCF-101 and HMDB51, leveraging robust pre-training on TikTok videos. The outcomes of this transfer learning process are presented in Table 3.
Dataset Top-1 (%) Top-5 (%)
Kinetics-400[4] 81.1 94.8
Something-Something V2[3] 69.2 91.6
Table 2: Direct Fine-Tuning Results With TikTok Videos Pre-training
Dataset Top-1 (%) Top-5 (%)
HMDB-51[5] 83.9 96.7
UCF-101[7] 97.5 99.6
@@
The fine-tuning results highlight the significant potential of using large, unlabeled video collections from platforms like TikTok for advancing human action recognition tasks. This approach capitalizes on the rich, varied nature of content available on social media, which, when used for pre-training, enhances the model’s ability to generalize across different datasets and recognition scenarios. The robust performance across both direct and transfer learning paths underscores the value of TikTok videos in training models to tackle complex action recognition challenges.
5 Extrapolating the Required Pre-Training Dataset Size
In Figure LABEL:fig:accplot, we find a monotone increasing accuracy score versus data volume. In the spirit of Floares et al. [2], we attempt to extrapolate the number of preprocessed videos for a desired accuracy by curve fitting to the number of data points trained so far. Because the accuracy maxes out at 100%, we choose to fit curves that are monotonically increasing with a horizontal asymptote at $11$. This makes any CDF of a random variable a viable choice for curve fitting. By iterating over the random variables in Scipy [8], we choose the top 10 distributions that fit the accuracy curve. Table 4 displays the number of pre-training data points required to achieve high accuracy levels (between 95% and 99%) according to the inverse CDF for common statistical models.
We validated these projections by pre-training a model with 9157 data points and fine-tuning the model using the same methods applied in the initial projections. The outcome of this validation was notably consistent with our predictive model. The model achieved an accuracy of 94.93%, closely aligning with the expected performance as indicated by the inverse CDF projections for this dataset size.
This validation not only reinforces the reliability of our statistical model but also highlights the practicality of using such predictive techniques in determining dataset size requirements. The fact that a dataset size of just 9157 points could yield an accuracy so close to 95% – a benchmark often targeted in action recognition tasks – is a testament to the efficiency of our approach. These results suggest that the requirement for massive datasets in the pre-training phase may be reconsidered, especially when computational resources are limited. This could mark a significant shift in how datasets are compiled and utilized in machine learning, particularly in fields where data acquisition is a challenge.
Table 4: Number of pre-training data points required to achieve high accuracy levels (between 95% and 99%) according to the inverse CDF for common statistical models.
Model Desired Accuracy
95% 96% 97% 98% 99%
t 9157 12388 18323 31892 82574
Burr XII 9090 12209 17857 30511 76240
InvGamma 9072 12166 17750 30204 74859
BetaPrime 9072 12165 17748 30200 74843
F 9072 12166 17749 30203 74855
GenPareto 9072 12164 17744 30184 74759
Lomax 9072 12164 17744 30184 74759
LogLaplace 9072 12164 17744 30184 74760
Pareto 9072 12164 17743 30183 74755
InvWeibull 9043 12101 17596 29772 72915

### v4

@@ Abstract
The increasing variety and quantity of tagged multimedia content on a variety of online platforms offer a unique opportunity to advance the field of human action recognition. In this study, we utilize 283,582 unique, unlabeled TikTok video clips, categorized into 386 hashtags, to train a domain-specific foundation model for action recognition. We employ VideoMAE V2, an advanced model integrating Masked Autoencoders (MAE) with Vision Transformers (ViT), pre-trained on this diverse collection of unstructured videos. Our model, fine-tuned on established action recognition benchmarks such as UCF101 and HMDB51, achieves state-of-the-art results: 99.05% on UCF101, 86.08% on HMDB51, 85.51% on Kinetics-400, and 74.27% on Something-Something V2 using the ViT-giant backbone. These results highlight the potential of using unstructured and unlabeled videos as a valuable source of diverse and dynamic content for training foundation models. Our investigation confirms that while initial increases in pre-training data volume significantly enhance model performance, the gains diminish as the dataset size continues to expand. Our findings emphasize two critical axioms in self-supervised learning for computer vision: (1) additional pre-training data can yield diminishing benefits for some datasets and (2) quality is more important than quantity in self-supervised learning, especially when building foundation models.
keywords:
Self-Supervised Learning, Action Recognition, Vision Transformers, TikTok Video Pre-training
@@ 1 Introduction
Action recognition is a fundamental computer vision task that involves identifying and classifying human actions within video sequences. This technology is critically important in various applications, including security systems, interactive gaming, and healthcare monitoring. As social media platforms have grown, TikTok has become a prominent source of diverse human action videos. In response, we have developed a specialized foundation model trained on a carefully curated set of 283,582 TikTok video clips. This dataset, known as TikTokActions, represents a wide array of human activities and is designed to reflect real-world scenarios and cultural diversities. It includes unique, non-standard human actions, filling gaps in existing action recognition datasets and expanding the variety of human activities that current models can recognize.
@@
Contrasting with the Kinetics dataset[1], which also sources user-generated content but from YouTube, our TikTokActions dataset is composed exclusively of TikTok videos. These are characterized by their dynamic and spontaneous nature, often featuring rapid transitions and creative interactions that are less common in YouTube’s typically longer and more structured videos. This difference provides unique training challenges, enhancing the ability of models to interpret a broad spectrum of human actions in unpredictable environments.
@@
We conducted a series of rigorous experiments to evaluate the efficacy of the TikTokActions dataset for pre-training computer vision models on action-related tasks. Using the advanced VideoMAE V2 model equipped with ViT-base and ViT-giant backbones[2], we explored the dataset’s capabilities. The ViT-base provided a baseline for comparison against other models with similar architectures, while the ViT-giant was used to assess the maximum potential of our dataset. Subsequent fine-tuning on well-known datasets such as UCF101[3], HMDB51[4], Kinetics[1], and Something-Something V2[5] revealed that models pre-trained on TikTokActions often performed comparably to, or even better than, those trained solely on these traditional benchmarks or those pre-trained on significantly larger datasets, where enhancements tend to be incremental.
Additionally, we investigated the relationship between the number of pre-training videos and model performance on downstream tasks. By training VideoMAE V2[2] on increasing subsets of our collection, from 1,000 videos up to 6,000, we observed that performance remained robust, suggesting that massive datasets are not always essential for effective pre-training. This discovery emphasizes the efficiency of using online, unlabeled videos as a source for pre-training, challenging the prevailing reliance on large-scale video datasets for optimal model performance. Prior to deploying such methodologies in practical applications, we urge the research community to consider the ethical implications of using publicly available data for training deep learning models, particularly in terms of data scraping and privacy concerns.
@@ 2.1 Human Action Recognition Datasets
The evolution of action recognition research has been significantly shaped by the development and availability of benchmark datasets, which guide and validate advancements in computer vision methodologies. The UCF101[3] dataset, known for its diverse array of human actions, has been instrumental in pioneering research within this domain. Likewise, the HMDB-51[4] dataset, despite its smaller scale, offers a rich variety of human actions captured under realistic conditions, providing essential challenges for model evaluation.
Introduced in 2017, the Kinetics Human Action Video dataset [1] became a cornerstone in this field with over 400 distinct human action classes derived from YouTube videos. It set a new benchmark, catalyzing further research and development in action recognition technologies. Similarly, AVA [6] provides densely labeled video segments that allow for action localization in addition to recognition, offering a deeper understanding of context and action dynamics.
Emerging datasets such as Ego4D [7], which focuses on first-person video analysis, and HowTo100M [8], which leverages instructional videos for action recognition, reflect the diversifying sources and modalities being explored. EPIC-Kitchens [9], ActivityNet [10], and THUMOS [11] offer further variations in complexity and scenarios, pushing the envelope on what models can understand and predict.
Recent datasets have addressed limitations of earlier collections. The Something-Something dataset [5] focuses on human-object interactions within everyday scenarios, offering a unique perspective that highlights the complexity of routine human activities. Moments in Time [12] and Multi-Moment Dataset [13] extend this by providing large-scale resources for temporal action understanding, capturing short videos that encompass a wide range of activities. The recent Tencent-MVSE dataset [14], another substantial contribution, is designed for multi-modal video similarity evaluation, adding another layer of complexity to action recognition tasks.
Building upon the diversity and scope of existing datasets, the TikTokActions dataset is meticulously designed to capture the rapidly evolving landscape of social media-driven human interactions. Unlike traditional datasets, which often focus on pre-defined, standard action categories, TikTokActions delves into a broader spectrum of spontaneous and culturally varied human behaviors, which are inherently more aligned with contemporary social media trends. This alignment is critical as platforms like TikTok foster unique user-generated content that is not only highly dynamic but also reflective of current societal norms and interactions.
In direct comparison to the Kinetics dataset, which largely aggregates content from YouTube, TikTokActions offers an array of actions driven by the unique constraints and creativity of TikTok’s platform, such as shorter video lengths and a more diverse global user base. This distinction is significant as it introduces new challenges in action recognition, particularly in recognizing quick, context-driven actions that are less prevalent in the typically longer-form content of YouTube.
To substantiate the dataset’s novelty and utility, we conducted a thorough analysis of the action categories present in TikTokActions compared to those in Kinetics and other major benchmarks. Our findings reveal that TikTokActions includes a higher proportion of non-standard actions, such as specific dance moves, challenge-based activities, and regional cultural expressions, which are underrepresented in other datasets. This inclusion not only enhances the dataset’s coverage of real-world scenarios but also provides a more granular understanding of human actions within the digitally connected world.
By providing detailed statistics on the distribution of these categories, along with examples in our supplementary materials, we demonstrate the unique value of TikTokActions in filling the existing gaps within the field of action recognition. This detailed comparison and analysis affirm that TikTokActions is not just an addition to the plethora of action recognition datasets but a necessary evolution to keep pace with the changing dynamics of human activity in the digital age.
@@ 2.2 Computer Vision Foundation Models
Foundation models in computer vision have revolutionized the way we approach diverse tasks across modalities, offering scalable solutions that extend from image to video understanding. The Florence model [15], Specifically, encapsulates the progression of foundation models, adeptly offering transitions from broad scene outlines to detailed object-centric views and from static imagery to dynamic video sequences. Additionally, its effectiveness spans various modalities, from RGB imagery to textual data, establishing new standards in tasks like classification, object detection, retrieval, and particularly in action recognition.
@@
Significant advancements have also been seen with the introduction of VideoMAE V2 [2]. This self-supervised pre-training methodology stands out for its ability to train on datasets exceeding 1 million videos. With its innovative dual masking strategy and progressive training approach, VideoMAE V2 sets new standards in action recognition. Empirical evaluations demonstrate its superior performance across multiple benchmarks, emphasizing its effectiveness in real-world scenarios.
Further developments include ViViT [16] and TimeSformer [17], which adapt transformer architectures specifically for video, offering improvements in handling the temporal dynamics of action sequences. These models demonstrate the potential of transformers in capturing complex, time-related patterns in video data.
The Quo Vadis model [18] introduced a new approach by integrating the Kinetics dataset to set a performance benchmark, inspiring subsequent models to focus on improving temporal understanding and action recognition capabilities. Similarly, Temporal Relational Reasoning in videos [19] provided insights into the importance of temporal dynamics, influencing how subsequent models like Non-local Neural Networks [20] handle long-range dependencies within video sequences.
A Closer Look at Spatiotemporal Convolutions [21] further explored the convolutional network architectures, focusing on how variations in spatial and temporal modules affect performance on action recognition tasks. This study paved the way for more specialized models like the Video Action Transformer Network [22], which introduced an approach to directly model relationships between actions and corresponding contextual features in videos.
Furthermore, 2D CNNs such as TSN [23], TSM [24], TANet [25], and TDN [26] have been influential by effectively incorporating temporal information, a critical aspect of video understanding.
The use of 3D CNNs such as I3D [18], R(2+1)D [21], ARTNet [27], and Slow-Fast [28] has further pushed the envelope in spatial-temporal feature integration, enhancing model performance on complex video tasks.
In the domain of transformers, advancements have continued with models such as MViT [29] and Video Swin Transformer [30] which have introduced ways to integrate spatial and temporal features more effectively, pushing the boundaries of action recognition performance. Additionally, UniFormer [31] integrates local and global interactions more efficiently, which is crucial for detailed action understanding in videos.
The diversity of foundation models extends to multi-modal approaches, where models like Perceiver IO [32] and Data2Vec [33] have been developed to handle various inputs ranging from text to video, showing flexibility and robustness in learning from heterogeneous data sources.
In the context of enhancing action recognition, SlowFast Networks [28] and X3D [34] introduce novel architectural choices that optimize the processing of temporal variations and the physical intensity of actions. These developments underscore the trend towards more specialized and efficient models that are not only capable of high performance but also adaptability across different action recognition scenarios.
Our TikTokActions dataset, with its focus on spontaneous and diverse user-generated content, provides a rich testing ground for these advanced models. By evaluating these models on our dataset, we aim to explore how well they can adapt to and interpret the nuanced, quick-transition actions typical of content found on modern social media platforms. The insights gained can help tailor these foundation models to better handle real-world variability and complexity in human actions.
In our efforts to push the frontiers of action recognition further, we also draw upon the extensive research documented in the Multi-dataset Training of Transformers [35], which explores the robustness gained from training across varied video datasets. Such approaches are crucial as they contribute to the generalization capabilities of models, preparing them for deployment in diverse, unstructured environments.
@@ 4 Performance on Benchmarks
In the domain of human action recognition, benchmarking performance is essential to understand the efficacy of a model. This section evaluates the performance of our model, initially pre-trained on the entire collection of 283,582 unlabeled TikTok video clips using VideoMAE V2[2]. We employed both the ViT-base and ViT-giant backbones for pre-training to assess the impact of model capacity. Subsequently, the model was fine-tuned on Kinetics-400 (K400). Following this, the model underwent further fine-tuning on UCF-101[3] and HMDB51[4], adopting a transfer learning approach akin to the experimental setups used in VideoMAE V2 studies. Additionally, we extended our evaluation to include the Something-Something V2[5] dataset, assessing the model’s adaptability and performance across varied action recognition contexts.
@@ 4.1 Pre-training on VideoMAE V2
VideoMAE V2[2] is a state-of-the-art masked autoencoding framework, notable for its flexibility in backbone selection, ranging from base to billion-level configurations. We opted for both the ViT-base and ViT-giant backbones. The ViT-base was used to provide a baseline comparison with other models using similarly sized backbones, while the ViT-giant was employed to explore the maximum potential of the dataset, aiming to achieve the highest possible accuracy. The generalization capability of VideoMAE V2 pre-trained ViTs[37] as video foundation models has been demonstrated across multiple downstream tasks, highlighting its robustness and adaptability.
Our training was conducted on 8 NVIDIA A100 40GB GPUs, utilizing the full set of 283,582 unlabeled TikTok videos. The pre-training on the ViT-base backbone, which ran for 800 epochs and lasted for 20 days, provided a comprehensive base for developing a robust model capable of understanding a wide array of human actions. The ViT-giant backbone was pre-trained under similar conditions, running for 1200 epochs over 30 days, to push the performance boundaries further.
@@ 4.2 Fine-Tuning on Established Datasets
After initial pre-training, the model underwent two distinct paths of fine-tuning. The first path involved direct fine-tuning on the Kinetics-400 and Something-Something V2 datasets to assess immediate applicability to broad action recognition tasks. The results of this direct fine-tuning are detailed in Table 2 and Table 3. The second path involved using the model fine-tuned on Kinetics-400 as an intermediary for transfer learning on UCF-101 and HMDB51, leveraging robust pre-training on TikTok videos. The outcomes of this transfer learning process are presented in Table 4.
Table 2: Results on the Kinetics-400 dataset. We scale the pre-training of VideoMAE V2 to billion-level ViT-g model with curated TikTok clips. We report the fine-tuning accuracy of multiple view fusion (5×3). Both models are pre-trained and fine-tuned at the input of 16×224×224 and sampling stride $\tau$ = 4.
method pre-train data data size epoch ViT-B ViT-g
MAE-ST [feichtenhofer2022masked] Kinetics400 0.24M 1600 81.3 -
VideoMAE V1 [38] Kinetics400 0.24M 1600 81.5 -
VideoMAE V2 [2] UnlabeledHybrid 1.35M 1200 81.5 87.2
VideoMAE V2 TikTok 0.28M 800 81.1 -
VideoMAE V2 TikTok 0.28M 1200 - 85.51
Table 3: Results on the Something-Something V2 dataset. We scale the pre-training of VideoMAE V2 to billion-level ViT-g model with curated TikTok clips. We report the fine-tuning accuracy of multiple view fusion (2×3). Both models are pre-trained and fine-tuned at the input of 16×224×224 and sampling stride $\tau$ = 2.
method pre-train data data size epoch ViT-B ViT-g
MAE-ST [feichtenhofer2022masked] Kinetics400 0.24M 1600 70.8 -
VideoMAE V1 [38] Sth-Sth V2 0.17M 2400 70.8 -
VideoMAE V2 [2] UnlabeledHybrid 1.35M 1200 71.2 77.0
VideoMAE V2 TikTok 0.28M 800 69.2 -
VideoMAE V2 TikTok 0.28M 1200 - 74.27
@@
Dataset ViT-b Top-1 (%) ViT-g Top-1 (%)
HMDB-51[4] 83.9 86.08
UCF-101[3] 97.5 99.05
Observations on data scaling. We studied the influence of pre-training data size on the performance of VideoMAE V2. In this experiment, we pre-trained the video models with ViT-B and ViT-g backbones on our TikTok dataset, which contains approximately 0.28M videos. The fine-tuning accuracy is detailed in Table 2 for Kinetics-400 and Table 3 for Something-Something V2. Despite using approximately five times fewer pre-training samples compared to the UnlabeledHybrid dataset (1.35M videos), our model achieved comparable results. Specifically, the ViT-B backbone pre-trained on TikTok data achieved 81.1% on Kinetics-400, which is only 0.4% lower than the 81.5% achieved using the much larger UnlabeledHybrid dataset. Similarly, for Something-Something V2, the ViT-B backbone pre-trained on TikTok data achieved 69.2%, compared to 71.2% with the UnlabeledHybrid dataset, showing a difference of 2%. These small differences highlight the efficacy of our foundation model on TikTok videos, which did not include any information from the original fine-tuned dataset.
Observations on model scaling. We also explored how different model capacities affect performance. The fine-tuning results for models pre-trained with ViT-B and ViT-g backbones are presented in Table 2 and Table 3. The ViT-g model, representing a billion-level parameter architecture, consistently demonstrated improved performance over the ViT-B model. For example, the ViT-g backbone pre-trained on TikTok data achieved 85.51% on Kinetics-400, compared to 81.1% for the ViT-B backbone, showing an improvement of 4.41%. Similarly, for Something-Something V2, compared to the ViT-B backbone, there is an improvement of 5.07%. These results suggest that increasing model capacity can further boost performance, validating the potential of large-scale pre-training on diverse video datasets like those from TikTok.
Observations on transfer learning. We further evaluated the performance of our pre-trained models by using Kinetics-400 fine-tuned models for transfer learning on HMDB-51 and UCF-101 datasets. The results are summarized in Table 4. The ViT-B model achieved 83.9% on HMDB-51 and 97.5% on UCF-101, while the ViT-g model achieved 86.08% on HMDB-51 and 99.05% on UCF-101.
@@
Overall, our findings highlight the significant potential of using large, unlabeled video collections from platforms like TikTok for advancing human action recognition tasks. This approach capitalizes on the rich, varied nature of content available on social media, which, when used for pre-training, enhances the model’s ability to generalize across different datasets and action recognition scenarios. The robust performance observed across both direct fine-tuning and transfer learning paths underscores the value of TikTok videos in training models for complex action recognition challenges. The TikTok dataset, which is unlabeled and lacks ground truth annotations, still produced promising results, demonstrating the model’s capacity to learn effectively from unstructured video content.
5 Dataset Size vs. Fine-Tuning Efficacy
To investigate the impact of pre-training data size on downstream model performance, we varied the size of the pre-training data from 1,000 to 6,000 videos, randomly selected from the TikTokActions dataset. Each data size level was used to pre-train the model in three independent runs to measure variance. Following pre-training, these models were fine-tuned on the UCF-101[3] dataset. The primary metric for evaluating model performance was the accuracy observed on the test set of UCF-101, allowing us to draw conclusions about the relationship between pre-training data size and fine-tuning effectiveness.
From the result in Figure  3, the volume of pre-trained data from the TikTokActions dataset was incrementally increased, and we observed a positive correlation with the Top-1 accuracy, particularly notable as the volume grew from 1,000 to 3,000 videos. However, subsequent increases in data volume yielded a reduced rate of improvement in Top-1 accuracy, indicating diminishing returns despite the larger dataset sizes. The Top-5 accuracy continued to show slight improvements, but the incremental benefit also appeared to diminish as the dataset size expanded beyond 3,000 videos. These findings suggest an asymptotic behavior in the benefit derived from increasing the volume of pre-training data for our models.
Figure 3: 100 Epoch Finetune: Final top-1 and top-5 accuracy by amount of pre-training data. We demonstrate the impact of the volume of pre-trained TikTok dataset on the fine-tuning accuracy for both Top-1 and Top-5 metrics.
6 Dataset Size vs. Fine-Tuning Efficacy
We conducted further experiments to demonstrate the efficacy of the collection in pre-training models for action recognition. We used a targeted subset of 6,900 high-quality clips sourced from 73 TikTok hashtags. From Figure  4, we can see the pre-trained model achieved a top-1 accuracy of 94.35% and a top-5 accuracy of 99.35% when fine-tuning on the UCF101[3]. This is remarkably close to the performance of models pre-trained on the UCF101 dataset itself (8155 from 101 classes), which exhibited a top-1 accuracy of 94.50% and a top-5 accuracy of 99.64%. The comparable accuracies highlight the mined dataset’s quality and its potential to serve as a substantial pre-training resource, even with fewer and more focused training samples. The model trained from scratch, without the benefit of such a rich dataset, performed significantly poorer, reinforcing the value of pre-training on high-quality, domain-specific data. This outcome attests to the robustness and generalization capabilities of models pre-trained on the curated dataset.
Figure 4: We compare the top-1 and top-5 accuracy of models pre-trained on TikTok actions, pre-trained from scratch, and pre-trained on UCF101 when fine-tuned to UCF101.